# On Simultaneous Information and Energy Transmission through Quantum Channels

### v6

@@ I Introduction
How physical systems process information and exchange energy is crucial to gaining insights into the workings of our universe. For example, the connections between entropy, information, and thermodynamics form the cornerstone of statistical mechanics. The study of phase transitions that involve an interplay between free energy and entropy, or Maxwell’s demon paradox and its resolution via Landauer’s principle are important examples. “Information is physical" as stated by Landauer [1] implies that the information processing capability of a device is inseparable from its physical properties. Quantum information processing has extended these ideas to the quantum domain [2, 3]. However, there is very little work considering quantum communication under physical constraints like energy.
@@
Classical information theory has shown a fundamental tradeoff between transmitting information and energy simultaneously using the same patterned energy signal, defined over a set of symbols that have different energy values [4, 5]. To transmit maximal information, the capacity-achieving input distribution over symbols should be used, whereas to maximize energy transmission, the most energetic symbol should be used all the time. There is a tension in trying to do both with the same signal. Although such information-theoretic analysis has inspired numerous practical engineering (communication and energy) systems such as energy-harvesting communications or powerline communications in the energy grid [6], it also addresses the fundamental relationship among basic physical quantities such as energy, matter, and information [7].
@@
Figure 1: The effect of threshold energy on the quality of information about the source sent to the receiver: (a) The information is encoded in quantum states, which are sent through a physical channel. At the output, a detector captures the states and decodes the information. However, the detector is activated only after a certain threshold of energy $B$, carried by the output states. (b) The effect of an increase in $B$ on the decoded information is shown.
@@
Communication has been a central topic in quantum information science: Sending classical or quantum information through a quantum channel has been studied extensively [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 3, 2, 21, 22]. Quantum communication is often performed using photons to transmit information over quantum channels. But photons also carry energy proportional to their wavelength. Here, we ask whether there is an analogous tension between transmitting information and energy simultaneously over quantum channels, as there is over classical channels. In particular, we characterize systems that simultaneously meet two goals:
@@
A classical picture of energy-constrained classical communication rate per unit cost was developed by Verdu [23], where he introduced the notion of capacity per unit cost function, which is the capacity of a channel under a cost constraint on the input signal. He gave a simplified formula for the capacity of such a channel per unit cost when there is a free symbol. Extending the same formulations to classical-quantum channels  [^1] We will revisit these definitions later in the manuscript., some previous works consider sending classical information through quantum channels under a maximum energy constraint on the input. Notably [25, 26] gave a simplified formula for capacity per unit cost when there exists a zero-cost state.
Here we consider requiring minimum received power rather than constraining transmit power, and consider the entire tradeoff curve to define quantum analogues of previously studied classical capacity-power functions [4]. We give new results and proofs regarding their properties. In particular, we analytically prove that the capacity-power function is concave for single-use unassisted communication (i.e. without any pre-shared entanglement) for the classical-quantum channel and for private communication through any degradable classical-quantum channel. We further prove additivity for these channels, so we do not need regularized versions to calculate the capacity-power function. We show these properties also hold for all noiseless channels when we restrict the set of input states to be pure quantum states. For general channels, we find that the capacity-power function is piecewise concave. We give a visual proof for this supported by numerical simulations. Lastly, we show how entropy of random quantum states under an energy constraint can be related to noiseless capacity-power functions.
@@
This provides both engineering utility in understanding the limits of quantum communications and broader insight into the basic physical interplay between energy and information. Figure 1 is a schematic diagram of the protocol we consider. This work deals with the fundamental tradeoff between transmitting energy and transmitting information over a single noisy line. The idea of energy-constrained private and quantum capacity was introduced in [27]. Unlike these past works and traditional transmitter power constraints in information theory, where small transmitted power is desired, here large received power is desired. Except for works on reversible computing [28], the fact that matter/energy must go along with information does not seem to have been considered in information theory. Similarly, the information carried in power transmission seems not to have been considered in power engineering.
@@ II Background
To explain the information-energy transmission setting, let us review the classical protocol that maximizes channel capacity under the constraint that the expected energy of the output should be greater than a given threshold [4]. First, we introduce some notation.
@@
[…]
Before proceeding, let us discuss some examples of the classical capacity-power function and methods of calculation [4]. If we have a classical memoryless binary noiseless channel without any energy constraint, then the capacity is $1$ and that capacity is achieved by the input distribution $P(0)=1/2$ and $P(1)=1/2$.
@@
The protocol for sending classical information through a quantum channel encodes information in quantum states at the transmitter. After sending through the quantum channel, the receiver performs measurement on the states to get the best possible information about the classical input distribution; this is the accessible information. If the transmitter prepares $\rho_{x}$, where $x=0,1\ldots,m$ with probabilities $p_{0},p_{1},\ldots,p_{m}$. Therefore, the transmitter has the state $\rho=\sum_{x}p_{x}\rho_{x}$. The receiver makes measurements using positive operator-valued measure (POVM) elements $\{E_{y}\}=\{E_{1},E_{2},\ldots,E_{k}\}$. The accessible information is determined by the Holevo bound [21] for a noiseless channel. and is given by:
@@
$C^{Q}(\mathcal{N})=\sup_{n}\>\frac{1}{n}{C_{n}^{Q}(\mathcal{N})}.$ (4)
@@
$=\max_{\text{all }p_{i},\rho_{i}}\left[S\left(\mathcal{N}(\rho)\right)-\sum_{i}p_{i}S\left(\mathcal{N}(\rho_{i})\right)\right],$ (5)
@@ III Capacity-Power Function of a Quantum Channel
$C^{Q}_{1}(B)=\underset{\begin{subarray}{c}\{p_{x},\rho_{x}\}:\\ Tr(H\mathcal{N}(\rho))\geq B\end{subarray}}{\max}\chi=\underset{\begin{subarray}{c}\{p_{x},\rho_{x}\}:\\ Tr(H\mathcal{N}(\rho))\geq B\end{subarray}}{\max}\left[S\left(\mathcal{N}(\rho)\right)-\sum_{i}p_{i}S\left(\mathcal{N}(\rho_{i})\right)\right]$ (6)
@@
[…]
The Holevo quantity $\chi$ is a concave function in probabilities and a convex function in the signal states[3]. This makes the computation of power-capacity for any general quantum channel difficult. However, we can design quantum channels by fixing the input quantum states. This seems more convenient operationally as well, where the experimentalist will have access to a fixed set of quantum states and have the freedom to change the classical probability distribution involved. Such channels are known as classical-quantum channels.
@@ IV Properties of the Capacity-Power Function
$\left[\mathrm{~{}S}\left(\sigma^{k}\right)-\sum_{x}p^{k}_{x}\mathrm{~{}S}\left(\sigma_{k}^{i}\right)\right]=C_{1}^{Q}(B_{k})\mbox{.}$ (10)
@@
$Tr(H^{\otimes 2}\mathcal{N}^{\otimes 2}(\rho^{\otimes 2}))=Tr(H\mathcal{N}(\rho_{1}))+Tr(H\mathcal{N}(\rho_{1}))\geq 2B,$ (22)
@@ IV.1.1 A Noiseless Quantum Channel
Let us take an example from Preskill [21] and add energy restrictions to it. Let us take
@@
Figure 11: 1.(a) Capacity-power function maximized over input probabilities for $|\alpha|=1,|\beta|=\sqrt{2},|\gamma|=\sqrt{3}$ for different values of noise parameter $p_{b}$ using (6). 1.(b) Variation of capacity-power function for all values of the noise parameter $p_{b}$. 1.(c) Two-dimension projection of capacity-power function in $p_{b}-B$ plane. 2.(a) Variation of capacity-power function $C_{1}^{Q}(B)$ for the depolarizing channel maximized over the ensembles of three random pure states with corresponding probabilities for different values of noise parameter $\lambda$. 2.(b) Variation of the capacity-power function for the depolarizing channel over all noise parameters $\lambda$. 2.(c) Two-dimensional projection of capacity-power function for the depolarizing channel.
@@
at each use of the channel. We know that [29]:
@@
The capacity-power function for this channel with three states in an ensemble for different values of noise parameter $p_{b}$ is shown in Fig. 11(1.(a)–1.(c)). There are protocols (one such one is given in [30]) to distinguish between coherent states using particular encoding and decoding schemes but all of them are bounded above by our results.
@@ V Capacity-Power Function and Random Quantum States
For an ensemble with pure input states that are orthogonal rank 1 projectors $\zeta=\{p_{i},\pi_{i}\}$, the noiseless channel capacity is given by the von Neumann entropy of the input ensemble, which is also the Shannon entropy of the probabilities. From random matrix theory we know if we take any uniformly distributed random pure state |$\ket{\psi}=\sum_{i}\sqrt{p_{i}}e^{i\phi_{i}}\ket{i}$, then the joint distribution of the $\{p_{i}\}$ will be uniform over the $N-1$ dimensional probability space [31]. To get this uniform distribution we need to draw each $p_{i}$ from a distribution
@@
The normalization condition of the $\{p_{i}\}$, i.e. that $\sum p_{i}=1$, is automatically satisfied for large $N$ as $\langle\sum p_{i}\rangle=1$ and $var(\sum p_{i})\propto N^{-1}$. So a typical random state from such an ensemble will have an entropy [31, 32]:
@@
The constant $0.422784$ is $1-\gamma$, where $\gamma$ is the Euler-Mascheroni constant. We give another way to arrive at the above expression in the appendix which essentially was first given in [33]. The variance of $\mathcal{H}$ goes to zero for large $N$ (see Fig. 13), akin to typicality arguments in information theory that rely on the concentration of measure. The entropy $\mathcal{H}$ for a typical random state only differs from the maximum possible entropy $\ln N$ by a constant. Hence, we do not need a maximization process to compute capacity corresponding to a large Hilbert space, since any randomly picked state has the same typical entropy. Therefore, capacity of a noiseless channel corresponding to a large dimension $N$ attains the value given by 35. This is also the channel capacity associated with typical states with no energy constraint.
@@
we can get the desired uniformity over the sliced space of probabilities [31] that satisfy $\sum_{i}p_{i}b_{i}=B$. The typical entropy associated with these conditions is
@@ VI Private Capacity-Power of a Quantum Channel
The private capacity of a quantum channel is the maximum rate at which classical information can be sent reliably and privately; operational definitions and coding theorems are given in prior work [27, 15, 34]. The protocol is shown in Fig. 14.
@@ VI.1 Private Capacity-Power Function of a Quantum Channel
$P_{1}^{Q}(\mathcal{N})=\max_{\{p_{i},\rho_{i}\}}\left(\chi_{AB}^{(1)}-\chi_{AE}^{(1)}\right)$ (49)
@@
[…]
@@
[…]
@@
[…]

### v7

@@ I Introduction
How physical systems process information and exchange energy is crucial to gaining insights into the workings of our universe. For example, the connections between entropy, information, and thermodynamics form the cornerstone of statistical mechanics. The study of phase transitions that involve an interplay between free energy and entropy, or Maxwell’s demon paradox and its resolution via Landauer’s principle are important examples. “Information is physical" as stated by Landauer launder1991 implies that the information processing capability of a device is inseparable from its physical properties. Quantum information processing has extended these ideas to the quantum domain nielsenchuang ; wilde_book . However, there is very little work considering quantum communication under physical constraints like energy.
The importance of simultaneous transmission of energy and information, two fundamental physical quantities, or the transmission of information under any constraint is of central importance in quantum information theory.
For example, quantum metrology deals with maximizing Fisher information in the presence of losses and under constraints of finite laser power and mean photon number. For example, let pure states be used in the arms of the interferometer and $\varphi$ be the parameter to be estimated. Then $\varphi$ dependent changes in the state in the interferometer is generated by the Hamiltonian,
$H_{\rm probe}=\varphi H.$ (1)
The quantum Cramér-Rao bound bound tells us that the measurement uncertainty is bounded from below as
$\delta\varphi\geq\frac{1}{2\sqrt{\langle\Delta^{2}H\rangle}},$ (2)
where $\langle\Delta^{2}H\rangle$ is the variance of $H$ caves_braunstein . It turns out that the mean photon number of the state in a wave-based interferometer with additional constraint like the requirement that the states in the two modes in the interferometer be both Gaussian constrains the minimum achievable uncertainty and therefore maximizes the Fisher information.
Rather than considering quantum metrology under constraints, here we consider quantum communication under power requirements. There are quantities like the mutual information, as fundamental as the Fisher information, which capture the rate of communication between two parties. In this work, we are interested in maximizing channel capacity under constraints, which is intimately related to mutual information.
@@
Classical information theory has shown a fundamental tradeoff between transmitting information and energy simultaneously using the same patterned energy signal, defined over a set of symbols that have different energy values varshney2008transporting ; grover2010shannon . To transmit maximal information, the capacity-achieving input distribution over symbols should be used, whereas to maximize energy transmission, the most energetic symbol should be used all the time. There is a tension in trying to do both with the same signal. Although such information-theoretic analysis has inspired numerous practical engineering (communication and energy) systems such as energy-harvesting communications or powerline communications in the energy grid clerckx2021wireless , it also addresses the fundamental relationship among basic physical quantities such as energy, matter, and information varshney2012 .
@@
Figure 1: Graphical summary: The effect of threshold energy on the information at the receiver: (a) The information is encoded in quantum states, which are sent through a physical channel. At the output, a detector captures the states and decodes the information. However, the detector is activated only after a certain threshold of energy $B$, carried by the output states. (b) A pictorial representation of the effect of an increase in $B$ on the decoded information. As the threshold of energy $B$ increases, information content at the receiver becomes poorer as shown by the increasingly mangled message at the receiver.
@@
Communication has been a central topic in quantum information science: Sending classical or quantum information through a quantum channel has been studied extensively Holevo2002 ; Holevo_some_estimates ; holevo2004entanglement ; Caves_bosonic_RevModPhys.66.481 ; gyongyosi2018survey ; Lyod_PhysRevA.55.1613 ; Shor_2003 ; Devetak2005 ; Additivity_of_the_classical_capacity_Shor ; Schaumacher ; HSW ; holevo_gen_signal ; Evaluating_capacities_of_bosonic_Gaussian_channels ; wilde_book ; nielsenchuang ; preskilllecture ; hastings2009superadditivity . Quantum communication is often performed using photons to transmit information over quantum channels. But photons also carry energy proportional to their wavelength. Here, we ask whether there is an analogous tension between transmitting information and energy simultaneously over quantum channels, as there is over classical channels. In particular, we characterize systems that simultaneously meet two goals:
@@
A classical picture of energy-constrained classical communication rate per unit cost was developed by Verdu verdu , where he introduced the notion of capacity per unit cost function, which is the capacity of a channel under a cost constraint on the input signal. He gave a simplified formula for the capacity of such a channel per unit cost when there is a free symbol. Extending the same formulations to classical-quantum channels  [^1] We will revisit these definitions later in the manuscript., some previous works consider sending classical information through quantum channels under a maximum energy constraint on the input. Notably wildecapacitycost ; PhysRevA.96.032340 gave a simplified formula for capacity per unit cost when there exists a zero-cost state.
Here we consider requiring minimum received power rather than constraining transmit power, and consider the entire tradeoff curve to define quantum analogues of previously studied classical capacity-power functions varshney2008transporting . We give new results and proofs regarding their properties. In particular, we analytically prove that the capacity-power function is concave for single-use unassisted communication (i.e. without any pre-shared entanglement) for the classical-quantum channel and for private communication through any degradable classical-quantum channel. We further prove additivity for these channels, so we do not need regularized versions to calculate the capacity-power function. We show these properties also hold for all noiseless channels when we restrict the set of input states to be pure quantum states. For general channels, we find that the capacity-power function is piecewise concave. We give a visual proof for this supported by numerical simulations. Lastly, we show how entropy of random quantum states under an energy constraint can be related to noiseless capacity-power functions.
@@
This provides both engineering utility in understanding the limits of quantum communications and broader insight into the basic physical interplay between energy and information. Figure 1 is a schematic diagram of the protocol we consider. This work deals with the fundamental tradeoff between transmitting energy and transmitting information over a single noisy line. The idea of energy-constrained private and quantum capacity was introduced in energy_constrained_private_and_quantum_capacity . Unlike these past works and traditional transmitter power constraints in information theory, where small transmitted power is desired, here large received power is desired. Except for works on reversible computing Varshney2010 , the fact that matter/energy must go along with information does not seem to have been considered in information theory. Similarly, the information carried in power transmission seems not to have been considered in power engineering.
@@ II Background
To explain the information-energy transmission setting, let us review the classical protocol that maximizes channel capacity under the constraint that the expected energy of the output should be greater than a given threshold varshney2008transporting . First, we introduce some notation.
@@
$C(B)=\underset{n\rightarrow\infty}{\lim}\frac{1}{n}C_{n}(B)\mbox{.}$
Before proceeding, let us discuss some examples of the classical capacity-power function and methods of calculation varshney2008transporting . If we have a classical memoryless binary noiseless channel without any energy constraint, then the capacity is $1$ and that capacity is achieved by the input distribution $P(0)=1/2$ and $P(1)=1/2$.
@@
The protocol for sending classical information through a quantum channel encodes information in quantum states at the transmitter. After sending through the quantum channel, the receiver performs measurement on the states to get the best possible information about the classical input distribution; this is the accessible information. If the transmitter prepares $\rho_{x}$, where $x=0,1\ldots,m$ with probabilities $p_{0},p_{1},\ldots,p_{m}$. Therefore, the transmitter has the state $\rho=\sum_{x}p_{x}\rho_{x}$. The receiver makes measurements using positive operator-valued measure (POVM) elements $\{E_{y}\}=\{E_{1},E_{2},\ldots,E_{k}\}$. The accessible information is determined by the Holevo bound preskilllecture for a noiseless channel.and is given by:
@@
$C^{Q}(\mathcal{N})=\underset{n\rightarrow\infty}{\lim}\>\frac{1}{n}{C_{n}^{Q}(\mathcal{N})}.$ (6)
@@
$=\max_{\text{all }\{p_{x},\rho_{x}\}}\left[S\left(\mathcal{N}(\rho)\right)-\sum_{x}p_{x}S\left(\mathcal{N}(\rho_{x})\right)\right]$ (7)
@@ III Capacity-Power Function of a Quantum Channel
$C^{Q}_{1}(B)=\underset{\begin{subarray}{c}\{p_{x},\rho_{x}\}:\\ Tr(H\mathcal{N}(\rho))\geq B\end{subarray}}{\max}\chi=\underset{\begin{subarray}{c}\{p_{x},\rho_{x}\}:\\ Tr(H\mathcal{N}(\rho))\geq B\end{subarray}}{\max}\left[S\left(\mathcal{N}(\rho)\right)-\sum_{x}p_{x}S\left(\mathcal{N}(\rho_{x})\right)\right]$ (8)
@@
$C^{Q}(B)=\underset{n\rightarrow\infty}{\lim}\>\frac{1}{n}C^{Q}_{n}(B)\mbox{.}$ (10)
The Holevo quantity $\chi$ is a concave function in probabilities and a convex function in the signal states wilde_book . This makes the computation of power-capacity for any general quantum channel difficult. However, we can design quantum channels by fixing the input quantum states. This seems more convenient operationally as well, where the experimentalist will have access to a fixed set of quantum states and have the freedom to change the classical probability distribution involved. Such channels are known as classical-quantum channels.
@@ IV Properties of the Capacity-Power Function
$\left[\mathrm{~{}S}\left(\sigma^{k}\right)-\sum_{x}p^{k}_{x}\mathrm{~{}S}\left(\sigma_{x}^{k}\right)\right]=C_{1}^{Q}(B_{k})\mbox{.}$ (12)
@@
$Tr(H^{\otimes 2}\mathcal{N}^{\otimes 2}(\rho^{\otimes 2}))=Tr(H\mathcal{N}(\rho_{1}))+Tr(H\mathcal{N}(\rho_{2}))\geq 2B,$ (24)
@@
In the discussion above, we have considered input states that are separable. In the case of an entangled state, the channel capacity can be superadditive hastings2009superadditivity . Though not the central focus of our paper, we briefly discuss a specific case of entangled states with entanglement-breaking channels and study the conditions (on the channel $\mathcal{N}$ and the input states) under which the additivity property for capacity-power function holds. [See Supplemental Material at URL will be inserted by publisher]
@@ IV.1.1 A Noiseless Quantum Channel
Let us take an example from Preskill preskilllecture and add energy restrictions to it. Let us take
@@
at each use of the channel. We know that gerry_knight_2004 :
@@
The capacity-power function for this channel with three states in an ensemble for different values of noise parameter $p_{b}$ is shown in Fig. 12(1.(a)–1.(c)). There are protocols (one such one is given in sidhu2021unambiguous ) to distinguish between coherent states using particular encoding and decoding schemes but all of them are bounded above by our results.
Figure 12: 1.(a) Capacity-power function maximized over input probabilities for $\alpha=1,\beta=\sqrt{2}e^{2\pi i/3},\gamma=\sqrt{3}e^{4\pi i/3}$ for different values of noise parameter $p_{b}$ using (8). 1.(b) Variation of capacity-power function for all values of the noise parameter $p_{b}$. 1.(c) Two-dimension projection of capacity-power function in $p_{b}-B$ plane. 2.(a) Variation of capacity-power function $C_{1}^{Q}(B)$ for the depolarizing channel maximized over the probabilities for different values of noise parameter $\lambda$. We took our signal states to be fixed at |$\ket{\psi_{1}}=\ket{0}$, |$\ket{\psi_{2}}=\cos(\pi/3)\ket{0}+\sin(\pi/3)\ket{1}$ and |$\ket{\psi_{3}}=\cos(2\pi/3)\ket{0}+\sin(2\pi/3)\ket{1}$ 2.(b) Variation of the capacity-power function for the depolarizing channel over all noise parameters $\lambda$. 2.(c) Two-dimensional projection of capacity-power function for the depolarizing channel.
@@ V Capacity-Power Function and Random Quantum States
For an ensemble with pure input states that are orthogonal rank 1 projectors $\zeta=\{p_{i},\pi_{i}\}$, the noiseless channel capacity is given by the von Neumann entropy of the input ensemble, which is also the Shannon entropy of the probabilities. From random matrix theory we know if we take any uniformly distributed random pure state |$\ket{\psi}=\sum_{i}\sqrt{p_{i}}e^{i\phi_{i}}\ket{i}$, then the joint distribution of the $\{p_{i}\}$ will be uniform over the $N-1$ dimensional probability space wooters . To get this uniform distribution we need to draw each $p_{i}$ from a distribution
@@
The normalization condition of the $\{p_{i}\}$, i.e. that $\sum p_{i}=1$, is automatically satisfied for large $N$ as $\langle\sum p_{i}\rangle=1$ and $var(\sum p_{i})\propto N^{-1}$. So a typical random state from such an ensemble will have an entropy wooters ; Zyczkowski_1990 :
@@
The constant $0.422784$ is $1-\gamma$, where $\gamma$ is the Euler-Mascheroni constant. We give another way to arrive at the above expression [See Supplemental Material at URL will be inserted by publisher] which essentially was first given in madhok2008 . The variance of $\mathcal{H}$ goes to zero for large $N$ (see Fig. 13), akin to typicality arguments in information theory that rely on the concentration of measure. The entropy $\mathcal{H}$ for a typical random state only differs from the maximum possible entropy $\ln N$ by a constant. Hence, we do not need a maximization process to compute capacity corresponding to a large Hilbert space, since any randomly picked state has the same typical entropy. Therefore, capacity of a noiseless channel corresponding to a large dimension $N$ attains the value given by 37. This is also the channel capacity associated with typical states with no energy constraint.
@@
we can get the desired uniformity over the sliced space of probabilities wooters that satisfy $\sum_{i}p_{i}b_{i}=B$. The typical entropy associated with these conditions is
@@ VI Private Capacity-Power of a Quantum Channel
The private capacity of a quantum channel is the maximum rate at which classical information can be sent reliably and privately; operational definitions and coding theorems are given in prior work energy_constrained_private_and_quantum_capacity ; Devetak2005 ; Devetak_private . The protocol is shown in Fig. 14. Along with channel capacity, private capacity has been extensively studied wilde_book . It is natural to extend how this quantity behaves under energy constraints and properties of private capacity-power functions. As the reliability of information transmission is a highly desirable property, one may ask the question: How does reliable transfer of information behave when the goal is to also transmit energy to the receiver as well?
@@ VI.1 Private Capacity-Power Function of a Quantum Channel
$P_{1}^{Q}(\mathcal{N})=\max_{\{p_{x},\rho_{x}\}}\left(\chi_{AB}^{(1)}-\chi_{AE}^{(1)}\right)$ (51)
@@
$P^{Q}(\mathcal{N})=\underset{n\rightarrow\infty}{\lim}\frac{1}{n}P_{n}^{Q}(\mathcal{N}^{\otimes n}).$ (52)
@@
$P^{Q}(\mathcal{N})=\underset{n\rightarrow\infty}{\lim}\>\frac{1}{n}\max_{\{p_{x},\rho_{x}\}}\left(\chi_{AB}^{(n)}-\chi_{AE}^{(n)}\right).$ (53)
@@
$P^{Q}(B)=\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\left[\underset{\begin{subarray}{c}\{p_{X^{n}},\rho_{X^{n}}\}:\\ Tr(H^{\otimes n}\mathcal{N}^{\otimes n}(\rho^{\otimes n}))\geq nB\end{subarray}}{\max}\left(\chi_{AB}^{(n)}-\chi_{AE}^{(n)}\right)\right]$ (54)